A liker walk 53.1 degree north of east for 2.5km then due east for 2km. What is her total displacement from her starting point if you measure the distance along straight line?


1
Expert's answer
2021-11-10T10:08:44-0500

"x =2.5\times cos(53.1) + 2=3.5 m \\

y = 2.5\times sin(53.1)=2 m"

displacement = "sqrt(2^2+3.5^2)=4.03 m"